This morning the longitudinal driving of PR marionette has been balanced, then the noise injection has been repeated.
Fig 1, 2, 3: PR marionette angular motion transmitted by the longitudinal noise, before and after the balancing.
An attempt to engage the reallocation has been performed at low noise 1, at the first relock after the maintenance (fig 1). The ITF survived for 20 seconds, then an oscillation at 12 Hz growed. It could be the roll resonance of the payload (to be checked). Moreover, the behaviour at 0.75 Hz was not good: this should be a problem of stability at the crossing frequency.
The blending filters have to be improved, and some commissioning time is needed in order to test accurately the performance.